# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ Abstract.
The Feder-Vardi dichotomy conjecture for Constraint Satisfaction Problems (CSPs) with finite templates, confirmed independently by Bulatov and Zhuk, has a counterpart for infinite templates due to Bodirsky and Pinsker which remains wide open. We resolve several meta-problems on the scope of their conjecture.
Our first two main results provide two fundamental simplifications of this scope, one of structural, and the other one of algebraic nature. The former simplification implies that the conjecture is equivalent to its restriction to templates without algebraicity, a crucial assumption in the most powerful classification methods. The latter yields that the higher-arity invariants of any template within its scope can be assumed to be essentially injective, and hence any algebraic condition characterizing any complexity class within the conjecture must be satisfiable by injections, thus lifting the mystery of the better applicability of certain conditions over others.
Our third main result uses the first one to show that any tractable template within the scope serves, up to a Datalog-computable modification of it, as the witness of the tractability of a finite-domain Promise Constraint Satisfaction Problem (PCSP) by the so-called sandwiching method. This provides a strong hitherto unknown connection between infinite-domain CSPs and finite-domain PCSPs.
Key words and phrases:
(Promise) Constraint Satisfaction Problem; dichotomy conjecture; polymorphism; identity; algebraicity; homogeneity; finite boundedness
Michael Pinsker, Jakub Rydval, and Moritz Schöbi: This research was funded in whole or in part by the Austrian Science Fund (FWF) [I 5948]. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. Michael Pinsker and Christoph Spiess: This research is funded by the European Union (ERC, POCOCOP, 101071674). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ 1.1. The finite-domain CSP dichotomy theorem
Consider the concrete example of 3-COLORING. This prominent NP-complete problem can be modeled as $\mathrm{CSP}(\mathfrak{K}_{3})$, where $\mathfrak{K}_{3}$ stands for the complete graph on three vertices: indeed, a homomorphism to $\mathfrak{K}_{3}$ from a given structure $\mathfrak{J}$ in the same signature (i.e. from a given graph $\mathfrak{J}$) is simply a mapping that does not map any two adjacent vertices of $\mathfrak{J}$ to the same vertex of $\mathfrak{K}_{3}$. A key observation in the theory of CSPs is that the NP-completeness of $\mathrm{CSP}(\mathfrak{K}_{3})$ can be extended to CSPs of other structures which can “simulate” $\mathfrak{K}_{3}$ over their domain or over a finite power thereof using primitive positive (pp) formulas. Enhancing this by the additional observation that homomorphically equivalent templates have equal CSPs, one arrives at the notion of pp-constructibility [9]. If $\mathfrak{A^{\prime}}$ is pp-constructible from $\mathfrak{A}$, then $\mathrm{CSP}(\mathfrak{A^{\prime}})$ is reducible to $\mathrm{CSP}(\mathfrak{A})$ in logarithmic space; this uniform reduction between CSP templates is sufficiently powerful, as we will see below, to describe all NP-hardness amongst finite-domain CSPs.
In the early 2000s, the field of finite-domain constraint satisfaction quickly rose in fame, in particular due to the discovery of tight connections with universal algebra. The most basic of these connections is that two structures with identical domains have the same sets of pp-definable relations if and only if they have the same sets of polymorphisms [53, 52]. Here, a polymorphism of a relational structure $\mathfrak{A}$ is simply a homomorphism from a finite power of $\mathfrak{A}$ into $\mathfrak{A}$ itself; we denote by $\mathrm{Pol}(\mathfrak{A})$ the set of all polymorphisms of $\mathfrak{A}$, the polymorphism clone of $\mathfrak{A}$. Since taking expansions by primitive positive relations does not lead to an increase in complexity, the study of finite-domain CSPs is subsumed by the study of finite algebras. Over the past two decades, the link between universal algebra and CSPs was gradually refined. After an intermediate stop at pp-interpretations [37], today we know that also pp-constructibility between finite relational structures is fully encoded in their polymorphism clones [9, Theorem 1.3]: $\mathfrak{A^{\prime}}$ is pp-constructible from $\mathfrak{A}$ if and only if $\mathrm{Pol}(\mathfrak{A^{\prime}})$ satisfies all height 1 identities of $\mathrm{Pol}(\mathfrak{A})$.
@@
Let $\mathfrak{A}$ be a finite relational structure. Then either $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ and $\mathrm{CSP}(\mathfrak{A})$ is NP-complete, or $\mathrm{Pol}(\mathfrak{A})$ satisfies the Siggers identity $s(x,y,z,x,y,z)\approx s(y,z,x,z,x,y)$ and $\mathrm{CSP}(\mathfrak{A})$ is solvable in polynomial time.
@@ 1.2. The infinite-domain CSP tractability conjecture
The quest for a CSP framework that includes such problems whilst staying within NP and allowing for an algebraic approach akin to the one for finite templates started with the popularization of $\omega$-categoricity by Bodirsky and Nešetřil [29] as a sufficient structural restriction ensuring the latter: for $\omega$-categorical structures $\mathfrak{A}$, pp-definability is determined by their polymorphism clone and, as was subsequently shown, so are the general reduction of pp-interpretability [31] as well as the pp-constructibility of $\mathfrak{K}_{3}$ [9]. The property of $\omega$-categoricity of $\mathfrak{A}$ can be interpreted as $\mathfrak{A}$ being “finite modulo automorphisms”: more precisely, the action of its automorphism group $\mathrm{Aut}(\mathfrak{A})$ on $d$-tuples has only finitely many orbits for all finite $d\geq 1$. It is, however, a mathematical property with little computational bearing: the CSPs of $\omega$-categorical structures can be monstrous from this perspective [49, 48]. For this reason, and based on strong empirical evidence, Bodirsky and Pinsker [33] identified a proper subclass of $\omega$-categorical structures as a candidate for an algebraic complexity dichotomy extending the theorem of Bulatov and Zhuk which does not seem to suffer from this deficiency.
Their first requirement on $\mathfrak{A}$ is that the orbit under $\mathrm{Aut}(\mathfrak{A})$ of any $d$-tuple be determined by the relations that hold on it: we call a relational structure $\mathfrak{A}$ homogeneous if every isomorphism between its finite substructures extends to an automorphism of $\mathfrak{A}$. This gives an effective way of representing orbits, which are central to $\omega$-categoricity. They then additionally impose an effective way of determining whether a given finite substructure is contained in $\mathfrak{A}$, which places the CSP in NP: a relational structure $\mathfrak{A}$ over a finite signature $\tau$ is finitely bounded [56] if there exists a finite set $\mathcal{N}$ of finite $\tau$-structures (bounds) such that a finite $\tau$-structure embeds into $\mathfrak{A}$ if and only if it does not embed any member of $\mathcal{N}$; equivalently, the finite substructures are up to isomorphism precisely the finite models of some universal first-order sentence [16, Lemma 2.3.14] (see also [27, 65] for more details).
@@
A standard example of a finitely bounded homogeneous structure is $(\mathbb{Q};<)$. Its finite substructures are all finite strict linear orders, which can be axiomatized by irreflexivity, totality, and transitivity:
$\forall x,y,z\;\;\big{(}\;(\neg x<x)\;\wedge\;(x<y\vee y<x)\;\wedge\;(x<y\wedge y<z\Rightarrow x<z)\;\big{)}.$
Every isomorphism between two such finite substructures can be extended to an automorphism of $(\mathbb{Q};<)$, e.g. by a piecewise affine transformation.
Finite boundedness and homogeneity taken together are very strong an assumption. However, Bodirsky and Pinsker observed that for all practical purposes, in particular the applicability of polymorphisms to complexity as well as containment in NP, it is sufficient that the template $\mathfrak{A}$ be a first-order reduct of a structure $\mathfrak{B}$ enjoying these, i.e. first-order definable therein. This yields sufficient flexibility to model a huge class of computational problems which includes in particular the ones mentioned above. In fact, requiring $\mathfrak{A}$ to be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$, i.e. obtained from $\mathfrak{B}$ by forgetting relations, turns out to be equivalent [3, Proposition 7], and we shall find this approach convenient.
@@
The CSP of the template $(\mathbb{Q};\operatorname{Betw})$, first-order defined by $\operatorname{Betw}(x,y,z):\Leftrightarrow(x<y<z)\vee(z<y<x)$ in the structure $(\mathbb{Q};<)$, is the classical NP-complete betweenness problem from the 1970s [62].
The following formulation of the Bodirsky-Pinsker conjecture takes into account later progress [9, 10, 11, 7].
Conjecture 1.4 (Bodirsky and Pinsker 2011 [33]).
Let $\mathfrak{A}$ be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. Then one of the following holds.
@@
  • $\mathrm{Pol}(\mathfrak{A})$ satisfies the pseudo-Siggers identity $\alpha\circ s(x,y,z,x,y,z)\approx\beta\circ s(y,z,x,z,x,y)$, and $\mathrm{CSP}(\mathfrak{A})$ is polynomial-time solvable.
@@
2. Three meta-questions: our contributions
The present article aims to resolve three meta-questions around Conjecture 1.4, of which the first concerns its scope; the second the algebraic invariants of templates therein; and the third its connection with the rapidly evolving field of (finite-domain) Promise Constraint Satisfaction Problems. More precisely, we consider the following questions:
@@ 2. Three meta-questions: our contributions
    Are there significant additional structural assumptions that can be imposed onto the structures of Conjecture 1.4 without loss of generality, i.e. without affecting the truth of the conjecture?
@@
    Are there algorithmic connections between CSPs from Conjecture 1.4 and Promise Constraint Satisfaction Problems?
@@ 2.1. Algebraicity is irrelevant
One such assumption is that the template $\mathfrak{A}$, or even the structure $\mathfrak{B}$ in which it is first-order defined, has no algebraicity. It is present virtually everywhere in the infinite-domain CSP literature [25]; and in particular an important prerequisite in most of the general results of the theory of smooth approximations developed by Mottet and Pinsker [61], as well as in all results of Bodirsky and Greiner [20, 21] about CSPs of generic superpositions of theories. We say that a structure $\mathfrak{A}$ has no algebraicity (in the group-theoretic sense) if, for every $k\geq 1$ and every tuple $\bar{a}\in A^{k}$, the automorphisms of $\mathfrak{A}$ which stabilize $\bar{a}$ do not stabilize any $a^{\prime}\notin\bar{a}$. In $\omega$-categorical structures this is the case if and only if, for every $k\geq 1$, every tuple $\bar{a}\in A^{k}$, and every $a^{\prime}\notin\bar{a}$, it is not possible to first-order define the unary relation $\{a^{\prime}\}$ using $\bar{a}$ as parameters; in other words, any non-trivial first-order property relative to $\bar{a}$ is always satisfied by infinitely many elements, which implies that some arguments become easier compared to finite structures (whereas naturally, many other arguments become harder). We will show that the following strengthening of this property (for model-complete cores, see Proposition 4.1) can be assumed when resolving Conjecture 1.4: a structure $\mathfrak{A}$ is CSP-injective if every finite structure that homomorphically maps to $\mathfrak{A}$ also does so injectively; in other words, if there is a solution to an instance of $\mathrm{CSP}(\mathfrak{A})$, then there is also an injective one. CSP-injectivity played an essential role in the universal-algebraic proof of the complexity dichotomy for Monotone Monadic SNP [25] and, more recently, certain finitely bounded homogeneous uniform hypergraphs [60].
@@
The templates $(\mathbb{Q};<)$ and $(\mathbb{Q};\operatorname{Betw})$ are clearly CSP-injective and have no algebraicity, whereas the equivalence relation $(V;\sim)$ with infinitely many classes of size 3 is not CSP-injective and has algebraicity.
In our theorem enforcing CSP-injectivity, one of the properties preserved by our construction is that of the structure $\mathfrak{B}$ being Ramsey. Although of central importance in classification methods [32, 63], its precise definition is not essential to the present paper and therefore omitted; we refer the reader to [51] for details about structural Ramsey theory.
Theorem 2.2.
Let $\mathfrak{A}$ be a non-trivial reduct of a countable structure $\mathfrak{B}$ over a finite relational signature. Then there exists a CSP-injective reduct $\mathbb{Q}^{\mathfrak{A}}$ of a countably infinite structure without algebraicity $(\mathbb{Q};<)^{\mathfrak{B}}$ over a finite relational signature such that $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ and $\mathrm{CSP}(\mathfrak{A})$ are Datalog-interreducible. Moreover:
@@
    If $\mathfrak{B}$ is $\omega$-categorical, then $(\mathbb{Q};<)^{\mathfrak{B}}$ is $\omega$-categorical as well. In this case, $\mathbb{Q}^{\mathfrak{A}}$ is a model-complete core if and only if $\mathfrak{A}$ is a model-complete core.
@@
    If $\mathfrak{B}$ is homogeneous, then $(\mathbb{Q};<)^{\mathfrak{B}}$ is homogeneous as well. In this case, $(\mathbb{Q};<)^{\mathfrak{B}}$ is Ramsey if and only if $\mathfrak{B}$ is Ramsey.
@@
    If $\mathfrak{B}$ is finitely bounded homogeneous, then $(\mathbb{Q};<)^{\mathfrak{B}}$ is finitely bounded homogeneous as well.
@@
    If the number of orbits of $d$-tuples of $\mathfrak{A}$ has growth slower than $2^{2^{d}}$, then $\mathbb{Q}^{\mathfrak{A}}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$.
Let us remark that it is well-known and not hard to see that reducts of homogeneous structures in a finite relational signature have less than double exponential orbit growth [56], and hence item 4 of Theorem 2.2 applies to the scope of Conjecture 1.4.

### v2

# Three Fundamental Questions in Modern Infinite-Domain Constraint Satisfaction

@@ Abstract.
The Feder-Vardi dichotomy conjecture for Constraint Satisfaction Problems (CSPs) with finite templates, confirmed independently by Bulatov and Zhuk, has an extension to certain well-behaved infinite templates due to Bodirsky and Pinsker which remains wide open. We provide answers to three fundamental questions on the scope of the Bodirsky-Pinsker conjecture.
Our first two main results provide two simplifications of this scope, one of structural, and the other one of algebraic nature. The former simplification implies that the conjecture is equivalent to its restriction to templates without algebraicity, a crucial assumption in the most powerful classification methods. The latter yields that the higher-arity invariants of any template within its scope can be assumed to be essentially injective, and any algebraic condition characterizing any complexity class within the conjecture closed under Datalog reductions must be satisfiable by injections, thus lifting the mystery of the better applicability of certain conditions over others.
Our third main result uses the first one to show that any non-trivially tractable template within the scope serves, up to a Datalog-computable modification of it, as the witness of the tractability of a non-finitely tractable finite-domain Promise Constraint Satisfaction Problem (PCSP) by the so-called sandwich method. This generalizes a recent result of Mottet and provides a strong hitherto unknown connection between the Bodirsky-Pinsker conjecture and finite-domain PCSPs.
@@ 1.1. The finite-domain CSP dichotomy theorem
Consider the concrete example of 3-COLORING. This prominent NP-complete problem can be modeled as $\mathrm{CSP}(\mathfrak{K}_{3})$, where $\mathfrak{K}_{3}$ stands for the complete graph on three vertices: indeed, a homomorphism to $\mathfrak{K}_{3}$ from a given structure $\mathfrak{J}$ in the same signature (i.e. from a given graph $\mathfrak{J}$) is simply a mapping that does not map any two adjacent vertices of $\mathfrak{J}$ to the same vertex of $\mathfrak{K}_{3}$. A key observation in the theory of CSPs is that the NP-completeness of $\mathrm{CSP}(\mathfrak{K}_{3})$ can be extended to CSPs of other structures which can “simulate” $\mathfrak{K}_{3}$ over their domain or over a finite power thereof using primitive positive (pp) formulas. Enhancing this by the additional observation that homomorphically equivalent templates have equal CSPs, one arrives at the notion of pp-constructibility [9]. If $\mathfrak{A^{\prime}}$ is pp-constructible from $\mathfrak{A}$, then $\mathrm{CSP}(\mathfrak{A^{\prime}})$ is reducible to $\mathrm{CSP}(\mathfrak{A})$ in logarithmic space; moreover, this reduction can be formulated in the logic programming language Datalog. This uniform reduction between CSP templates is sufficiently powerful, as we will see below, to describe all NP-hardness amongst finite-domain CSPs.
In the early 2000s, the field of finite-domain constraint satisfaction quickly rose in fame, in particular due to the discovery of tight connections with universal algebra. The most basic of these connections is that two structures with identical domains have the same sets of pp-definable relations if and only if they have the same sets of polymorphisms [51, 50]. Here, a polymorphism of a relational structure $\mathfrak{A}$ is simply a homomorphism from a finite power of $\mathfrak{A}$ into $\mathfrak{A}$ itself; we denote by $\mathrm{Pol}(\mathfrak{A})$ the set of all polymorphisms of $\mathfrak{A}$, the polymorphism clone of $\mathfrak{A}$. Since taking expansions by primitive positive relations does not lead to an increase in complexity, the study of finite-domain CSPs is subsumed by the study of finite algebras. Over the past two decades, the link between universal algebra and CSPs was gradually refined. After an intermediate stop at pp-interpretations [36], today we know that also pp-constructibility between finite relational structures is fully encoded in their polymorphism clones [9, Theorem 1.3]: $\mathfrak{A^{\prime}}$ is pp-constructible from $\mathfrak{A}$ if and only if $\mathrm{Pol}(\mathfrak{A^{\prime}})$ satisfies all height-1 identities of $\mathrm{Pol}(\mathfrak{A})$.
@@
Let $\mathfrak{A}$ be a finite relational structure. Then either $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ and $\mathrm{CSP}(\mathfrak{A})$ is NP-complete, or $\mathrm{Pol}(\mathfrak{A})$ satisfies the Siggers identity
$s(x,y,z,x,y,z)\approx s(y,z,x,z,x,y)$
and $\mathrm{CSP}(\mathfrak{A})$ is solvable in polynomial time.
@@ 1.2. The infinite-domain CSP tractability conjecture
The quest for a CSP framework that includes such problems whilst staying within NP and allowing for an algebraic approach akin to the one for finite templates started with the popularization of $\omega$-categoricity by Bodirsky and Nešetřil [28] as a sufficient structural restriction ensuring the latter: for countable $\omega$-categorical structures, pp-definability is determined by their polymorphism clone and, as was subsequently shown, so are the general reduction of pp-interpretability [30] as well as the pp-constructibility of $\mathfrak{K}_{3}$ [9]. The $\omega$-categoricity of $\mathfrak{A}$ can be interpreted as $\mathfrak{A}$ being “finite modulo automorphisms”: more precisely, the action of its automorphism group $\mathrm{Aut}(\mathfrak{A})$ on $d$-tuples has only finitely many orbits for all finite $d\geq 1$. It is, however, a mathematical property with little computational bearing: the CSPs of $\omega$-categorical structures can be monstrous from this perspective [48, 47]. For this reason, and based on strong empirical evidence, Bodirsky and Pinsker [32] identified a proper subclass of countable $\omega$-categorical structures as a candidate for an algebraic complexity dichotomy extending the theorem of Bulatov and Zhuk which does not seem to suffer from this deficiency.
Their first requirement on $\mathfrak{A}$ is that the orbit under $\mathrm{Aut}(\mathfrak{A})$ of any $d$-tuple be determined by the relations that hold on it: we call a relational structure $\mathfrak{A}$ homogeneous if every isomorphism between its finite substructures extends to an automorphism of $\mathfrak{A}$. This gives an effective way of representing orbits, which are central to $\omega$-categoricity. They then additionally impose an effective way of determining whether a given finite substructure is contained in $\mathfrak{A}$, which places the CSP in NP: a relational structure $\mathfrak{A}$ over a finite signature $\tau$ is finitely bounded [54] if there exists a finite set $\mathcal{N}$ of finite $\tau$-structures (bounds) such that a finite $\tau$-structure embeds into $\mathfrak{A}$ if and only if it does not embed any member of $\mathcal{N}$; equivalently, the finite substructures are up to isomorphism precisely the finite models of some universal first-order sentence [15, Lemma 2.3.14] (see also [26, 63] for more details).
@@
A standard example of a countable finitely bounded homogeneous structure is $(\mathbb{Q};<)$. Its finite substructures are all finite strict linear orders, which can be axiomatized by irreflexivity, totality, and transitivity:
$\forall x,y,z\big{(}\neg(x<x)\wedge(x<y\vee x=y\vee y<x)\wedge(x<y\wedge y<z\Rightarrow x<z)\big{)}.$
Morevoer, every isomorphism between two such finite substructures can be extended to an automorphism of $(\mathbb{Q};<)$, e.g. by a piecewise affine transformation.
Finite boundedness and homogeneity taken together are very strong an assumption. Bodirsky and Pinsker observed that for all practical purposes, in particular the applicability of polymorphisms to complexity as well as containment in NP, it is sufficient that the template $\mathfrak{A}$ be a first-order reduct of a structure $\mathfrak{B}$ enjoying these, i.e. first-order definable therein. This yields sufficient flexibility to model a huge class of computational problems which includes in particular the ones mentioned above. In fact, requiring $\mathfrak{A}$ to be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$, i.e. obtained from $\mathfrak{B}$ by forgetting relations, turns out to be equivalent [3, Proposition 7], and we shall find this approach convenient.
@@
The CSP of the template $(\mathbb{Q};\operatorname{Betw})$, first-order defined by
$\operatorname{Betw}(x,y,z):\Leftrightarrow(x<y<z)\vee(z<y<x)$
in the structure $(\mathbb{Q};<)$, is the classical NP-complete betweenness problem from the 1970s [60].
The following formulation of the conjecture of Bodirsky and Pinsker takes into account later progress [9, 10, 11, 7].
Conjecture 1.4 (Bodirsky and Pinsker 2012 [32]).
Let $\mathfrak{A}$ be a reduct of a countable finitely bounded homogeneous structure $\mathfrak{B}$. Then exactly one of the following holds.
@@
  • $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$; in this case $\mathrm{Pol}(\mathfrak{A})$ satisfies the pseudo-Siggers identity
    $\alpha\circ s(x,y,z,x,y,z)\approx\beta\circ s(y,z,x,z,x,y)$
    and $\mathrm{CSP}(\mathfrak{A})$ is polynomial-time solvable.
@@
2. Three fundamental questions: our contributions
The present paper aims to answer three fundamental questions around Conjecture 1.4, of which the first concerns its scope; the second the algebraic invariants of templates therein; and the third its connection with the rapidly evolving field of (finite-domain) Promise Constraint Satisfaction Problems. More precisely, we consider the following questions:
@@ 2. Three fundamental questions: our contributions
    Are there significant additional structural assumptions, perhaps of model-theoretic nature, that can be imposed onto the structures of Conjecture 1.4 without loss of generality, i.e. without affecting the truth of the conjecture?
@@
    Are there algorithmic connections between CSPs from Conjecture 1.4 and finite-domain Promise Constraint Satisfaction Problems?
@@ 2.1. Algebraicity is irrelevant
One such assumption is that the template $\mathfrak{A}$, or even the structure $\mathfrak{B}$ in which it is first-order defined, has no algebraicity. It is present virtually everywhere in the infinite-domain CSP literature [24]; and in particular an important prerequisite in most of the general results of the theory of smooth approximations developed by Mottet and Pinsker [59], as well as in all results of Bodirsky and Greiner [19, 20] about CSPs of generic superpositions of theories. We say that a structure $\mathfrak{A}$ has no algebraicity (in the group-theoretic sense) if, for every $k\geq 1$ and every tuple $\bar{a}\in A^{k}$, the automorphisms of $\mathfrak{A}$ which stabilize $\bar{a}$ do not stabilize any $a^{\prime}\notin\bar{a}$. In $\omega$-categorical structures this is the case if and only if, for every $k\geq 1$, every tuple $\bar{a}\in A^{k}$, and every $a^{\prime}\notin\bar{a}$, it is not possible to first-order define the unary relation $\{a^{\prime}\}$ using $\bar{a}$ as parameters; in other words, any non-trivial first-order property relative to $\bar{a}$ is always satisfied by infinitely many elements, which implies that some arguments become easier compared to finite structures (whereas naturally, many other arguments become harder). We will show that the following strengthening of this property (for model-complete cores, see Proposition 4.1 in Section 4.1) can be assumed when resolving Conjecture 1.4: a structure $\mathfrak{A}$ is CSP-injective if every finite structure that homomorphically maps to $\mathfrak{A}$ also does so injectively; in other words, if there is a solution to an instance of $\mathrm{CSP}(\mathfrak{A})$, then there is also an injective one. CSP-injectivity played an essential role in the universal-algebraic proof of the complexity dichotomy for Monotone Monadic SNP [24] and, more recently, certain finitely bounded homogeneous uniform hypergraphs [58].
@@
The template $(\mathbb{Q};<)$ from Example 1.2 is CSP-injective because a digraph maps homomorphically to $(\mathbb{Q};<)$ if and only if it is acyclic; in this case there is always an injective homomorphism. It is a model-complete core, and hence has no algebraicity by Proposition 4.1 in Section 4.1. Its expansion $(\mathbb{Q};<)_{I_{4}}$ by the relation
$I_{4}\coloneqq\{(x,y,u,v)\mid x=y\Rightarrow u=v\}$
has the same automorphisms, and hence no algebraicity; it is no more CSP-injective by Propositions 4.2 and 4.3. The equivalence relation with countably many classes of size 2 has algebraicity: any automorphism stabilizing some element $a$ must also stabilize the other element of its class.
One of the properties preserved by our construction enforcing CSP-injectivity in Theorem 2.2 below is that of the structure $\mathfrak{B}$ being Ramsey. Although of central importance in classification methods [31, 61], its precise definition is not essential to the present paper and therefore omitted; we refer the reader to [49] for details about structural Ramsey theory.
Throughout the present paper, the notation $\mathfrak{A}\wr\mathfrak{B}$ stands for the relational wreath product of two structures $\mathfrak{A}$ and $\mathfrak{B}$, introduced formally in Section 4.2, and $\mathfrak{A}_{R}$ stands for the expansion of a structure $\mathfrak{A}$ by a relation $R$ defined over its domain.
Theorem 2.2 (No Algebraicity).
Let $\mathfrak{A}$ be a non-trivial reduct of a countable structure $\mathfrak{B}$ over a finite relational signature. Then $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is a non-trivial CSP-injective reduct of $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$, which is a countably infinite structure without algebraicity over a finite relational signature, and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ and $\mathrm{CSP}(\mathfrak{A})$ are Datalog-interreducible. Moreover:
@@
    If $\mathfrak{B}$ is $\omega$-categorical, then $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ is $\omega$-categorical as well. In this case:
    • $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is a model-complete core if and only if $\mathfrak{A}$ is a model-complete core.
@@
    If $\mathfrak{B}$ is homogeneous, then $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ is homogeneous as well. In this case:
    • $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ is Ramsey if and only if $\mathfrak{B}$ is Ramsey.
@@
    If $\mathfrak{B}$ is finitely bounded, then $\smash{\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}}$ is finitely bounded as well.
@@
    If the number of orbits of $d$-tuples of $\mathfrak{A}$ has growth slower than $2^{2^{d}}$, then:
    • $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$.
It is not hard to see that item 4 of Theorem 2.2 applies to reducts of homogeneous structures in a finite relational signature [54], in particular to all structures in the scope of Conjecture 1.4.